A Bibliography of Supercomputing ’2003

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 May 2011 Version 1.02

Abstract

This bibliography records articles presented at the Supercomputing ’2003 conference.

Title word cross-reference

10-Gigabit [cFJH^{+}03].

2 [ABH^{+}03]. 2003 [ACM03].

6 [MKFD03].

Accelerators [QFR03]. Access [WTR03]. Achieving [PKP03]. Adaptive [BC03]. adding [JTB^{+}03]. Algebraic [ABKP03]. Analysis [DL03b]. Applications [ABKP03]. Approach [FFP03].

Approximate [WZS03]. Architecture [SOB03]. Architectures [OCC^{+}03]. ASCI [PKP03]. Automatic [CMKH03]. Availability [NKB^{+}03]. Awareness [JTB^{+}03].

Balance [LC03]. Balancing [BC03]. Based [NKB^{+}03]. BCS [FFP03]. BCS-MPI [FFP03]. Beams [QFR03]. Beck [AW03]. billion [KTJT03]. Blocking [Nak03]. Bone [ABKP03]. Browsing [DHB^{+}03].

Cache [OCC^{+}03]. Cache-based [OCC^{+}03]. Cacheless [OCC^{+}03]. Case [cFJH^{+}03]. Catalog [SBC^{+}03]. Cell [QFR03]. Chaff [CW03]. Chaff-based [CW03]. CLI [Vog03]. CLI-based [Vog03]. Cluster [NKB^{+}03].

Cluster-Based [NKB^{+}03]. Clusters [TY03]. Coarse [DFA03]. Coarse-Grained [DFA03]. Code [KLK^{+}03]. Colliding [QFR03]. Communication [DL03b]. Communications [JUS03]. Comparison [LCW^{+}03]. Compiler [DL03b]. Computational [LD03]. Computations [WZS03]. Computer [TNO^{+}03]. Computers [ABB^{+}03]. Computing [Vog03]. Condors [BZH03]. Configurable [JUS03]. Conservative [YSF03]. Constraint [Kar03]. Contact [Nak03]. Contact/Impact [Kar03]. Contiguous [WTR03]. Cosmology [WFG03]. Counters [IWAK03]. Cray [DFIW03].

Daemon [MCSL03]. Data [TY03]. Databases [DHB^{+}03]. Decisions [YSF03]. Dedicated [WLNL03]. degrees [KTJT03]. Design [FFP03]. Detection [RHB03]. Dimensions [BBC03]. Disk [WB03]. Distributed [KLA^{+}03]. Domain [KLK^{+}03]. Domain-wise [KLK^{+}03]. Driven [CMKH03]. Dyn [WLNL03]. Dyn-MPI [WLNL03]. Dynamic [YSF03]. Dynamics [KC03].

Early [DFIW03]. Earth [Nak03]. Earthquake [MSB^{+}03]. Efficient [TY03]. Element [ABKP03]. Emulation [LC03]. Enabling [OSGR03]. Energy [QFR03]. Environment [KLA^{+}03]. Environments [YSF03]. Ethernet [cFJH^{+}03]. Evaluation [DFIW03]. Experience [ABH^{+}03]. Exploiting [DFA03]. Explorer [TNO^{+}03]. fast [YBZL03]. Fault [BCH^{+}03]. File [WB03]. Finite [ABKP03]. Flock [BZH03]. Fold [RHB03]. Forward [ABB^{+}03]. Framework [BC03]. freedom [KTJT03].

Galaxy [DPK^{+}03]. GAMESS [OSGR03]. GAMESS/DDI [OSGR03]. Generation [CMKH03]. Genomic [RHB03]. GeoFEM [Nak03]. Gigabit [cFJH^{+}03]. Grained [DFA03]. GRAPE [MKFD03]. GRAPE-6 [MKFD03]. Grid [DL03a]. Grid-Based [DPK^{+}03]. Grids [LD03]. GridSAT [CW03]. Gyrokinetic [KEL^{+}03].

Handling [WB03]. Hardware [JUS03]. Hardware-Based [CDPM03]. Heterogeneity [WB03]. Hierarchical [KC03]. High [NKB^{+}03]. High-Performance [NKB^{+}03]. Highly [BBC03]. HPC.NET [Vog03].

I/O [LkLC^{+}03]. IBM [ABKP03].

Identifying [MdSM^{+}03].

Igniting [ACM03]. Image [DHB^{+}03].

Implementations [LCW^{+}03]. implemented [KEL^{+}03]. Improve [YSF03]. Improvement [RHB03]. Improving [NKB^{+}03]. independent [YBZL03]. InfiniBand [LCW^{+}03]. Information [DL03a]. Innovation [ACM03]. Intelligent [JUS03]. Intensive [SBC^{+}03]. Interarrival [KC03]. Interface [LkLC^{+}03]. Internet [NKB^{+}03]. Interprocedural [DL03b]. Inverse [WZS03]. IPSAP [KLK^{+}03]. Irregular [BC03]. Iterative [Nak03].

Job [SOB03]. Jobs [JTB^{+}03].

kernel [YBZL03].

kernel-independent [YBZL03].

Languages [CMKH03]. Large [MSB^{+}03]. Large-Scale [MSB^{+}03]. Library [CMKH03]. Livny [AW03]. Load [LC03]. Location [TY03]. Logging [BCH^{+}03]. Logistical [DHB^{+}03].

Machines [Vog03]. Matrix [WZS03]. Mechanics [ABKP03]. Memory [IWAK03]. Merrimac [DHE^{+}03]. Mesh [Kar03]. Message [BCH^{+}03]. Metadata [SBC^{+}03]. method [YBZL03]. Micro [ABKP03]. Micro-Mechanics [ABKP03]. Million [RHB03]. Million-Fold [RHB03]. Missing [PKP03]. Model [OSGR03]. Modeling [ABB^{+}03]. Modular [JUS03]. Molecular [TNO^{+}03]. Monitoring [WR03]. Morphology [DPK^{+}03]. Movement [AW03]. MPI [LCW^{+}03]. MPICH [BCH^{+}03]. MPICH-V2 [BCH^{+}03]. MRNet [RAM03]. MTA [ABH^{+}03]. MTA-2 [ABH^{+}03]. Multi [Kar03]. Multi-Constraint [Kar03]. Multicast [RAM03]. Multicast/ Reduction [RAM03]. Multifrontal [KLK^{+}03]. Multigrid [ABKP03]. Multilevel [WZS03]. multipole [YBZL03]. Multiprocessor [WR03]. Myrinet [LCW^{+}03].

National [DPK^{+}03]. netCDF [LkLC^{+}03]. Network [RAM03]. Networking [DHB^{+}03]. Networks [cFJH^{+}03]. NIC [JUS03]. NIC-based [MFPP03]. Nodes [BCH^{+}03]. Non [WLNL03]. Non-Contiguous [WTR03]. Nondeterministic [DL03a]. Nonlinear [Nak03]. November [ACM03].

O [LkLC^{+}03]. Observatory [DPK^{+}03]. OpenMP [KKH03]. Operating [JTB^{+}03]. Optimal [PKP03]. Optimizing [KLA^{+}03]. organizing [TY03].

ParADE [KKH03].

Parallel [WZS03, LkLC^{+}03].

Parallelism [DFA03]. Particle [QFR03]. Particle-In-Cell [QFR03]. Partitioning [Kar03]. DDI [OSGR03]. Impact [Kar03]. Reduction [RAM03]. Performance [NKB^{+}03]. Pessimistic [BCH^{+}03]. Petaflops [TNO^{+}03]. Phoenix [ACM03]. Pipelined [DFA03]. Plank [AW03]. Plank-Beck [AW03]. Poisson [BBC03]. Preconditioners [WZS03]. Preconditioning [Nak03]. Predicted [YSF03]. Primitives [JUS03]. Problems [Nak03]. Processors [PKP03]. Profiling [IWAK03]. Programming [KKH03]. Programs [ABH^{+}03]. Protein [TNO^{+}03]. Protocol [TY03, MCSL03]. Purpose [TNO^{+}03].

Q [PKP03]. Quadrics [LCW^{+}03].

Quantifying [NKB^{+}03]. Queries [DL03a].

real [MKFD03]. Realistic [LD03].

Reduction [KLA^{+}03]. References [MdSM^{+}03]. Regularity [MdSM^{+}03]. Relational [DL03a]. Remote [DHB^{+}03]. Repeats [RHB03]. Resolution [ABB^{+}03].

SAT [CW03]. SC2003 [ACM03].

Scalability [JTB^{+}03]. Scalable [RAM03]. Scale [MSB^{+}03]. SCALLOP [BBC03]. Scheduling [YSF03]. Scientific [OCC^{+}03]. Selective [Nak03]. Self [TY03]. Self-organizing [TY03]. Sender [BCH^{+}03]. Service [DL03a]. Services [NKB^{+}03]. Shared [WB03]. Shared-Disk [WB03]. Simulation [QFR03]. Simulations [MSB^{+}03]. Simulator [Nak03]. SMP [OSGR03].

Software [RAM03]. Software-Based [RAM03]. Solver [BBC03]. Solvers [Nak03]. SP [ABKP03]. Space [WFG03]. Sparse [WZS03]. Spatial [MdSM^{+}03]. Special [TNO^{+}03]. Special-Purpose [TNO^{+}03]. Speed [RHB03]. Storage [TY03]. Streaming [KEL^{+}03]. Streams [DHE^{+}03]. Structural [KLK^{+}03]. Studies [AW03]. Study [cFJH^{+}03]. Suitable [Vog03]. Supercomputer [PKP03]. Supercomputing [DHE^{+}03]. Supernovae [WFG03]. Superscalar [OCC^{+}03]. Superscheduler [SOB03]. Support [DFA03]. Supporting [WLNL03]. Synthesizing [LD03]. System [TNO^{+}03]. Systems [WB03].

Technique [KLK^{+}03]. Telescoping [CMKH03]. terabyte [KTJT03]. teraflops [KTJT03]. Terascale [ABB^{+}03]. Tflops [MKFD03]. Three [BBC03]. Times [KC03]. Tolerant [BCH^{+}03]. Tool [MCSL03]. Tools [RAM03]. Toroidal [KEL^{+}03]. Traffic [LC03]. Traffic-based [LC03]. Trees [CDPM03]. tuning [MKFD03]. Type [CMKH03]. Type-Driven [CMKH03].

Unified [WR03]. Universe [WFG03]. Use [OSGR03]. Using [YSF03].

V2 [BCH^{+}03]. Variance [YSF03]. Vector [OCC^{+}03]. Very [MSB^{+}03]. Virtual [Vog03]. Visualization [DHB^{+}03]. Visualizing [MSB^{+}03]. Volatile [BCH^{+}03].

Whole [ABKP03]. wise [KLK^{+}03]. Workstations [cFJH^{+}03].

X1 [DFIW03].

References

Akcelik:2003:HRF

[ABB^{+}03] Volkan Akcelik, Jacobo Bielak, George Biros, Ioannis Epanomeritakis, Antonio Fernandez, Omar Ghattas, Eui Joong Kim, Julio Lopez, David O’Hallaron, Tiankai Tu, and John Urbanic. High resolution forward and inverse earthquake modeling on terascale computers. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                                   ????  URL http://www.sc-
                                                   conference.org/sc2003/inter_
                                                   cal/inter_cal_detail.php?eventid=
                                                   10707#0; http://www.sc-conference.
                                                   org/sc2003/paperpdfs/pap298.
                                                   pdf.

Anderson:2003:EES

[ABH^{+}03] Wendell Anderson, Preston Briggs, C. Stephen Hellberg, Daryl W. Hess, Alexei Khokhlov, Marco Lanzagorta, and Robert Rosenberg. Early experience with scientific programs on the Cray MTA-2. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

                                                   LCCN ???? URL http://www.
                                                   sc-conference.org/sc2003/inter_
                                                   cal/inter_cal_detail.php?eventid=
                                                   10717#2; http://www.sc-conference.
                                                   org/sc2003/paperpdfs/pap271.
                                                   pdf.

Adams:2003:AAM

[ABKP03] Mark F. Adams, Harun H. Bayraktar, Tony M. Keaveny, and Panayiotis Papadopoulos. Applications of algebraic multigrid to large-scale finite element analysis of whole bone micro-mechanics on the IBM SP. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10694#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap211.
pdf.

ACM:2003:SII

[ACM03] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

Allen:2003:LPB

[AW03] Matthew S. Allen and Rich Wolski. The Livny and Plank-Beck problems: Studies in data movement on the computational grid. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10718#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap266.
pdf.

Balls:2003:SHS

[BBC03] Gregory T. Balls, Scott B.

Baden, and Phillip Colella. SCALLOP: A highly scalable parallel Poisson solver in three dimensions. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

                                        LCCN ???? URL http://www.
                                        sc-conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10703#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap201.
                                        pdf.

Barker:2003:EFD

[BC03] Kevin J. Barker and Nikos P. Chrisochoides. An evaluation of a framework for the dynamic load balancing of highly adaptive and irregular parallel applications. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10708#2; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap268.
                                        pdf.

Bouteiller:2003:MVF

[BCH^{+}03] Aurelien Bouteiller, Franck Cappello, Thomas Herault, Geraud Krawezik, Pierre Lemarinier, and Frederic Magniette. MPICH-V2: a fault tolerant MPI for volatile nodes based on pessimistic sender based message logging. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10696#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap209.
                                        pdf.

Butt:2003:SOF

[BZH03] Ali Raza Butt, Rongmei Zhang, and Y. Charlie Hu. A selforganizing flock of Condors.

In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10690#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap265.
pdf.

Coll:2003:SHB

[CDPM03] Salvador Coll, Jose Duato, Fabrizio Petrini, and Francisco J. Mora. Scalable hardware-based multicast trees. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10702#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap300.
pdf.

Feng:2003:OGE

[cFJH^{+}03] Wu chun Feng, Justin, Hurwitz, Harvey B. Newman, Sylvain Ravot, Roger Les Cottrell, Olivier Martin, Fabrizio Coccetti, Cheng Jin, David Wei, and Steven Low. Optimizing 10-gigabit Ethernet in networks of workstations, clusters, and grids: A case study. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10702#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap293.
pdf.

Chauhan:2003:ATD

[CMKH03] Arun Chauhan, Cheryl McCosh, Ken Kennedy, and Richard Hanson. Automatic type-driven library generation for telescoping languages. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

                                        LCCN ???? URL http://www.
                                        sc-conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10692#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap296.
                                        pdf.

Chrabakh:2003:GCB

[CW03] Wahid Chrabakh and Rich Wolski. GridSAT: A chaff-based distributed SAT solver for the grid. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10710#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap252.
                                        pdf.

Du:2003:CSE

[DFA03] Wei Du, Renato Ferreira, and Gagan Agrawal. Compiler support for exploiting coarsegrained pipelined parallelism. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10692#2; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap133.
                                        pdf.

Dunigan:2003:EEC

[DFIW03] Thomas H. Dunigan, Jr., Mark R.

Fahey, James B. White III, and Patrick H. Worley. Early evaluation of the Cray X1. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10706#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap183.
pdf.

Ding:2003:RVB

[DHB^{+}03] Jin Ding, Jian Huang, Micah Beck, Shaotao Liu, Terry Moore, and Stephen Soltesz. Remote visualization by browsing image based databases with logistical networking. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10691#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap245.
pdf.

Dally:2003:MSS

[DHE^{+}03] William J. Dally, Patrick Hanrahan, Mattan Erez, Timothy J. Knight, Francois Labonte, Jung-Ho Ahn, Nuwan Jayasena, Ujval J. Kapasi, Abhishek Das, Jayanth Gummaraju, and Ian Buck. Merrimac: Supercomputing with streams. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=

                                        10717#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap246.
                                        pdf.

Dinda:2003:NQR

[DL03a] Peter Dinda and Dong Lu. Nondeterministic queries in a relational grid information service. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10695#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap146.
                                        pdf.

Ding:2003:CAI

[DL03b] Yonghua Ding and Zhiyuan Li. A compiler analysis of interprocedural data communication. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10692#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap137.
                                        pdf.

Deelman:2003:GBG

[DPK^{+}03] Ewa Deelman, Raymond Plante, Carl Kesselman, Gurmeet Singh, Mei Su, Gretchen Greene, Robert Hanisch, Niall Gaffney, Antonio Volpicelli, James Annis, Vijay Sekhri, Tamas Budavari, Maria Nieto-Santisteban, William O’Mullane, David Bohlender, Tom McGlynn, Arnold Rots, and Olga Pevunova. Grid-based galaxy morphology analysis for the National Virtual Observatory. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10718#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap282.
pdf.

Fernandez:2003:BMN

[FFP03] Juan Fernandez, Eitan Frachtenberg, and Fabrizio Petrini. BCS-MPI: a new approach in the system software design for large-scale parallel computers. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10716#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap306.
pdf.

Itzkowitz:2003:MPU

[IWAK03] Marty Itzkowitz, Brian J. N.

Wylie, Christopher Aoki, and Nicolai Kosche. Memory profiling using hardware counters. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10721#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap182.
pdf.

Jones:2003:ISP

[JTB^{+}03] Terry Jones, William Tuel, Larry Brenner, Jeff Fier, Patrick Caffrey, Shawn Dawson, Rob Neely, Robert Blackmore, Brian Maskell, Paul Tomlinson, and Mark Roberts. Improving the scalability of parallel jobs by adding parallel awareness to the operating system. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http://

                                        www.sc-conference.org/sc2003/
                                        inter_cal/inter_cal_detail.
                                        php?eventid=10716#0; http:/
                                        /www.sc-conference.org/sc2003/
                                        paperpdfs/pap136.pdf.

Jaganathan:2003:CNP

[JUS03] Ranjesh G. Jaganathan, Keith D. Underwood, and Ron R. Sass. A configurable network protocol for cluster based communications using modular hardware primitives on an intelligent NIC. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10702#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap199.
                                        pdf.

Karypis:2003:MCM

[Kar03] George Karypis. Multi-constraint mesh partitioning for contact/ impact computations. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http://

                                        www.sc-conference.org/sc2003/
                                        inter_cal/inter_cal_detail.
                                        php?eventid=10703#2; http:/
                                        /www.sc-conference.org/sc2003/
                                        paperpdfs/pap302.pdf.

Kleban:2003:HDI

[KC03] Stephen D. Kleban and Scott H.

Clearwater. Hierarchical dynamics, interarrival times, and performance. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10696#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap222.
pdf.

Klasky:2003:GBP

[KEL^{+}03] Scott Alan Klasky, Stephane Ethier, Zhihong Lin, Kevin Martins, Doug McCune, and Ravi Samtaney. Grid-based parallel data streaming implemented for the gyrokinetic toroidal code. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10722#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap207.
pdf.

Kee:2003:POP

[KKH03] Yang-Suk Kee, Jin-Soo Kim, and Soonhoi Ha. ParADE: An OpenMP programming environment for SMP cluster systems. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10708#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap130.
pdf.

Kurc:2003:ORC

[KLA^{+}03] Tahsin Kurc, Feng Lee, Gagan Agrawal, Umit Catalyurek, Renato Ferreira, and Joel Saltz. Optimizing reduction computations in a distributed environment. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10695#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap135.
pdf.

Kim:2003:IHP

[KLK^{+}03] Seung Jo Kim, Chang Sung Lee, Jeong Ho Kim, Minsu Joh, and Sangsan Lee. IPSAP : A high-performance parallel finite element code for largescale structural analysis based on domain-wise multifrontal technique. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10707#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap238.
pdf.

Komatitsch:2003:BDF

[KTJT03] Dimitri Komatitsch, Seiji Tsuboi, Chen Ji, and Jeroen Tromp. A 14.6 billion degrees of freedom, 5 teraflops, 2.5 terabyte earthquake simulation on the Earth Simulator. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10711#1; http://www.sc-conference.

org/sc2003/paperpdfs/pap124.
pdf.

Liu:2003:TBL

[LC03] Xin Liu and Andrew A. Chien.

Traffic-based load balance for scalable network emulation. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10690#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap260.
pdf.

Liu:2003:PCM

[LCW^{+}03] Jiuxing Liu, Balasubramanian Chandrasekaran, Jiesheng Wu, Weihang Jiang, Sushmitha Kini, Weikuan Yu, Darius Buntinas, Pete Wyckoff, and D. K. Panda. Performance comparison of MPI implementations over InfiniBand, Myrinet and Quadrics. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10696#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap310.
pdf.

Lu:2003:SRC

[LD03] Dong Lu and Peter August Dinda. Synthesizing realistic computational grids. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http://

www.sc-conference.org/sc2003/
inter_cal/inter_cal_detail.
php?eventid=10690#0; http:/

                                        /www.sc-conference.org/sc2003/
                                        paperpdfs/pap180.pdf.

Li:2003:PNH

[LkLC^{+}03] Jianwei Li, Wei keng Liao, Alok Choudhary, Robert Ross, Rajeev Thakur, William Gropp, Rob Latham, Andrew Siegel, Brad Gallagher, and Michael Zingale. Parallel netCDF: A highperformance scientific I/O interface. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10722#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap258.
                                        pdf.

Miller:2003:TDP

[MCSL03] Barton Miller, Ana Cortés, Miquel Senar, and Miron Livny. The Tool Daemon Protocol. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10687#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap187.
                                        pdf.

Mohan:2003:IES

[MdSM^{+}03] Tushar Mohan, Bronis R. de Supinski, Sally A. McKee, Frank Mueller, Andy Yoo, and Martin Schulz. Identifying and exploiting spatial regularity in data memory references. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-

conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10721#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap290.
pdf.

Moody:2003:SNB

[MFPP03] Adam Moody, Juan Fernandez, Fabrizio Petrini, and Dhabaleswar K. Panda. Scalable NIC-based reduction on largescale clusters. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10716#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap316.
pdf.

Makino:2003:PET

[MKFD03] Junichiro Makino, Eiichiro Kokubo, Toshiyuki Fukushige, and Hiroshi Daisaka. Performance evaluation and tuning of GRAPE-6 — towards 40 ‘real’ Tflops. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10711#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap116.
pdf.

Ma:2003:VVL

[MSB^{+}03] Kwan-Liu Ma, Aleksander Stompel, Jacobo Bielak, Omar Ghattas, and Eui Joong Kim. Visualizing very large-scale earthquake simulations. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.

                                        sc-conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10691#2; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap287.
                                        pdf.

Nakajima:2003:PIS

[Nak03] Kengo Nakajima. Parallel iterative solvers of GeoFEM with selective blocking preconditioning for nonlinear contact problems on the Earth Simulator. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10703#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap155.
                                        pdf.

Nagaraja:2003:QIA

[NKB^{+}03] Kiran Nagaraja, Neeraj Krishnan, Ricardo Bianchini, Richard P. Martin, and Thu D. Nguyen. Quantifying and improving the availability of high-performance cluster-based Internet services. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10686#2; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap213.
                                        pdf.

Oliker:2003:ECB

[OCC^{+}03] Leonid Oliker, Andrew Canning, Jonathan Carter, John Shalf, David Skinner, Stephane Ethier, Rupak Biswas, Jahed Djomehri, and Rob Van der Wijngaart.

Evaluation of cache-based superscalar and cacheless vector architectures for scientific computations. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10706#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap255.
pdf.

Olson:2003:EEU

[OSGR03] Ryan M. Olson, Michael W.

Schmidt, Mark S. Gordon, and Alistair P. Rendell. Enabling the efficient use of SMP clusters: The GAMESS/DDI model. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10691#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap263.
pdf.

Petrini:2003:CMS

[PKP03] Fabrizio Petrini, Darren J. Kerbyson, and Scott Pakin. The case of the missing supercomputer performance: Achieving optimal performance on the 8,192 processors of ASCI Q. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10706#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap301.
pdf.

Qiang:2003:PPC

[QFR03] Ji Qiang, Miguel A. Furman, and Robert D. Ryne. Parallel particle-in-cell simulation of colliding beams in high energy accelerators. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10694#2; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap223.
                                        pdf.

Roth:2003:MSB

[RAM03] Philip C. Roth, Dorian C. Arnold, and Barton P. Miller. MRNet: A software-based multicast/reduction network for scalable tools. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

                                        LCCN ???? URL http://www.
                                        sc-conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10687#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap192.
                                        pdf.

Romein:2003:MFS

[RHB03] John W. Romein, Jaap Heringa, and Henri E. Bal. A millionfold speed improvement in genomic repeats detection. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http://

                                        www.sc-conference.org/sc2003/
                                        inter_cal/inter_cal_detail.
                                        php?eventid=10710#0; http:/
                                        /www.sc-conference.org/sc2003/
                                        paperpdfs/pap189.pdf.

Singh:2003:MCS

[SBC^{+}03] Gurmeet Singh, Shishir Bharathi, Ann Chervenak, Ewa Deelman, Carl Kesselman, Mary Manohar, Sonal Patil, and Laura Pearlman. A metadata catalog service for data intensive applications. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10718#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap242.
pdf.

Shan:2003:JSA

[SOB03] Hongzhang Shan, Leonid Oliker, and Rupak Biswas. Job superscheduler architecture and performance in computational grid environments. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10695#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap267.
pdf.

Taiji:2003:PEP

[TNO^{+}03] Makoto Taiji, Tetsu Narumi, Yousuke Ohno, Noriyuki Futatsugi, Atsushi Suenaga, Naoki Takada, and Akihiko Konagaya. Protein Explorer: A petaflops special-purpose computer system for molecular dynamics simulations. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=

                                        10717#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap168.
                                        pdf.

Tang:2003:EDL

[TY03] Hong Tang and Tao Yang. An efficient data location protocol for self-organizing storage clusters. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10686#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap299.
                                        pdf.

Vogels:2003:HNC

[Vog03] Werner Vogels. HPC.NET — are CLI-based virtual machines suitable for high performance computing? In ACM [ACM03], page ?? ISBN 1-58113-695-1.

                                        LCCN ???? URL http://www.
                                        sc-conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10710#2; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap251.
                                        pdf.

Wu:2003:HHS

[WB03] Changxun Wu and Randal Burns. Handling heterogeneity in shared-disk file systems. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10686#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap132.
                                        pdf.

Warren:2003:SSM

[WFG03] Michael S. Warren, Chris L.

Fryer, and M. Patrick Goda. The Space Simulator: Modeling the universe from supernovae to cosmology. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10711#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap226.
pdf.

Weatherly:2003:DMS

[WLNL03] D. Brent Weatherly, David K.

Lowenthal, Mario Nakazawa, and Franklin Lowenthal. Dyn-MPI: Supporting MPI on non dedicated clusters. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10708#1; http://www.sc-conference.
org/sc2003/paperpdfs/pap126.
pdf.

Wisniewski:2003:EUS

[WR03] Robert W. Wisniewski and Bryan Rosenburg. Efficient, unified, and scalable performance monitoring for multiprocessor operating systems. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-
conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10721#0; http://www.sc-conference.
org/sc2003/paperpdfs/pap121.
pdf.

Worringen:2003:FPN

[WTR03] Joachim Worringen, Jesper Larson Traff, and Hubert Ritzdorf. Fast parallel non-contiguous file access. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

                                        LCCN ???? URL http://www.
                                        sc-conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10722#0; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap319.
                                        pdf.

Wang:2003:PMS

[WZS03] Kai Wang, Jun Zhang, and Chi Shen. Parallel multilevel sparse approximate inverse preconditioners in large sparse matrix computations. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.sc-
                                        conference.org/sc2003/inter_
                                        cal/inter_cal_detail.php?eventid=
                                        10694#1; http://www.sc-conference.
                                        org/sc2003/paperpdfs/pap109.
                                        pdf.

Ying:2003:NPK

[YBZL03] Lexing Ying, George Biros, Denis Zorin, and Harper Langston. A new parallel kernel-independent fast multipole method. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http://

                                        www.sc-conference.org/sc2003/
                                        inter_cal/inter_cal_detail.
                                        php?eventid=10707#2; http:/
                                        /www.sc-conference.org/sc2003/
                                        paperpdfs/pap166.pdf.

Yang:2003:CSU

[YSF03] Lingyun Yang, Jennifer M.

Schopf, and Ian Foster. Conservative scheduling: Using predicted variance to improve scheduling decisions in dynamic environments. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/inter_
cal/inter_cal_detail.php?eventid=
10687#2; http://www.sc-conference.
org/sc2003/paperpdfs/pap236.
pdf.